1.) A 238 mL solution has an osmotic pressure of 0.398 atm at a temperature of 32°C. If 0.312 g of solute was added to the solution, what is the molar mass of the solute?
1
Expert's answer
2020-05-14T14:34:17-0400

The osmotic pressure of the solution can be expressed as


"\pi=i\frac{nRT}{V},\\
n=\frac{\pi V}{RT}."

if the van't Hoff coefficient is "i=1".

The molar mass is


"\Mu=\frac{m}{n}=\frac{mRT}{\pi V}=\frac{0.312\cdot8.314(32+273)}{0.398\cdot10^5\cdot238\cdot10^{-6}}=83.5\text{ g/mol}."